9/8 / 3/5 of (3/4+3/5)
Answer by Cromlix(4381)

9/8 / 3/5 of 3/4 + 3/5
9 * 5/8 * 3 of 27/20
= 45/24 of 27/20
Cancels down to give you
= 9/8 of 9/4
= 81/32
= 2 17/32
Hope this helps
:-)